OpenStudy (anonymous):

What is the value of the expression (-2)^5 a -10 b -32 c 10 d 32

OpenStudy (anonymous):

i think d

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

\[(-2)^5=(-1)^5(-2)^5=-32\]

OpenStudy (jdoe0001):

\(\bf (-2)^5\implies -2\cdot -2\cdot -2\cdot -2\cdot -2\)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

u to

OpenStudy (anonymous):

Wait go 1 more question

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What is the value of the expression –5x^4 when x = –2? a -40 b -80 c 40 d 80

OpenStudy (jdoe0001):

\(\bf 5{\color{blue}{ x}}^4 \qquad x = -2\qquad \implies 5({\color{blue}{ -2}})^4\\ \quad \\ \implies 5\cdot -2\cdot -2\cdot -2\cdot -2\)

OpenStudy (anonymous):

what is ur name

OpenStudy (anonymous):

Mine

OpenStudy (anonymous):

what